# Evaluating Feynman Integrals through differential equations and series expansions

### v1

@@ 1 Introduction
The Standard Model (SM) is the theory that classifies all elementary particles and describes their interactions. During the years it has proven itself very successful in explaining and predicting with extreme precisions a big variety of phenomena, spanning several orders of magnitude. The discovery of the existence of the Higgs boson [1, 2], which has been confirmed in 2012 at the Large Hadron Collider (LHC), is one of its greater success. Despite its incredible achievements, however, the SM has some unsolved problems. For example, it does not include gravity in its description of reality, it is not able to explain dark matter and dark energy nor the matter-antimatter asymmetry. The presence of these problems hints that the Standard Model is not yet complete and an extension is needed. Nowadays there exist multiple theories which are suitable candidates for an extension of the SM, namely Grand Unified Theories (GUT), Supersymmetry, String Theory etc., but no clear evidence of physics beyond the SM has been found. Although during the high luminosity run at the LHC we may still be able to make new exciting discoveries, physicists have started to pursue another path to explore the unknown territory of new physics, that is looking for small deviations between theoretical predictions and experimental data. The presence of these small deviations could give, indeed, precious insight on how to extend the SM. For this reason theoretical physicists in the latest years have focused their efforts on obtaining predictions as accurate as possible. These calculations, however, are very far from trivial and in order to achieve a percent, or even sub-percent, level of accuracy one must have control on all the sources of theory uncertainties. These include, for instance, Parton Distributions Functions (PDFs), input scheme and higher order corrections, computed either in Quantum Chromodynamics (QCD) or in the Electro-Weak (EW) sector. One of the main bottleneck for the inclusion of higher order corrections is the calculation of the so-called Feynman integrals [3, 4]. Indeed, in state-of-the-art problems their number can grow up to thousands or hundred of thousands, and hence a general and algorithmic approach is necessary for their evaluation. Fortunately, not all of them are independent one from the other, and by exploiting integration-by-parts (IBPs) identities [5] one is able to find a set of independent integrals, such that all the others can be expressed in terms of these ones. These integrals are called Master Integrals (MIs). One of the most effective computational tools for evaluating the MIs is the method of differential equations, firstly proposed by Kotikov [6] and later improved by Remiddi and Gehrmann [7, 8, 9, 10]. The idea behind this method is that these integrals are functions of kinematics variables and internal masses, and by differentiating with respect to one of those variables one is able to obtain a first order linear differential equation whose unknowns are the MIs we are interested in. The problem, hence, shifts from integrating to solving a differential equation. The solution to these equations can be written in terms of one-dimensional iterated integrals, which, in many cases, correspond to some known classes of functions such as Harmonic Polylogarithms (HPLs) [11, 12, 13, 14, 15]. However, when increasing the number of loops, external legs or internal masses, an analytical solution in terms of known classes of functions can become extremely difficult to obtain. For this reason, in latest years the series expansion approach is gaining popularity, thanks also to many public implementations: AMFLow [16], DiffExp [17], Line [18] and SeaSyde [19]. Its main idea is to look for a series solution to the differential equations, so that the we are able to evaluate it easily at any point of the phase-space. A general difficulty that arises when calculating corrections at fixed perturbative order is the presence of internal unstable particles such as Ws and Zs. A complete description of resonances in perturbation theory requires a Dyson summation of self-energy insertions. In particular, it has been observed that the pole of the resummed propagator is a gauge-invariant complex quantity. Therefore it is possible to define the renormalized mass of unstable particles as the complex pole of the resummed propagator. This scheme is called complex mass scheme (CMS) [20, 21], and is necessary to correctly describe the kinematic region near the resonance. In order to perform a complete calculation in the CMS we need to be able to evaluate the master integrals also with complex internal masses, and the series expansion approach to the differential equation method can be exploited to this end.
@@ 2.2 Integration-by-parts identities
During state-of-the-art calculations, one usually has to compute up to hundred of thousands of FIs. However, not all of them are independent, but there exist linear relations among them. These relations go under the name of Integration-by-parts Identities (IBPs) and they follow from Gauss theorem in d-dimensions:
@@
where $v_{\mu}$ can be either a loop or an external momentum. The first step is to write down identities from Eq. 4, for different values of $v_{\mu}$ and for all the $\boldsymbol{\alpha}$ that appear in our calculation. Then we can solve the system by Gauss substitution rule, i.e. considering each equation, expressing one integral in terms of the others and substitute it in the remaining ones. This procedure can be approached in a systematic way by using Laporta algorithm [23]. This is done by assigning a weight to each integral and by systematically expressing integrals of higher weight in terms of others with lower weight. Laporta algorithm is extremely simple and elegant and it has been implemented in several public codes, such as Blade [24], Kira [25, 26], LiteRed [27], NeatIBP [28] and Reduze [29]. Its algebraic complexity, however, grows very fast with the number of loops, external legs and internal masses. For this reason, for state of the art problem, obtaining compact IBPs in a reasonable amount of time is one of the bottleneck of the calculations. After the implementation of IBPs we are able to express every integral, belonging to the a given integral family, in terms of a finite number of FIs which are called Master Integrals (MIs). These Master Integrals play the role of a basis in the space of Feynman integrals. Note that the term basis may be misleading, because a priori we do not know if this set of MIs is minimal, that is if there exist other relations between them and so if they are truly independent on from the other, anyway this term gives a good idea of what is going on.
@@ 3.2 Example: 1L QED vertex
A second comment regards the number of variables that appear in the system. A common procedure is to introduce adimensional variables, defined as ratios of kinematics invariants. This is done in order to reduce by one the total number of variables in the differential equations, thus simplifying the solution procedure. For example, one could introduce the variable $x=s/m^{2}$, so that the system in Eq. 13 becomes
@@ 3.3 Boundary Conditions
Before diving into how to solve the system of differential equations let us address the problem of finding the boundary conditions. The boundary conditions, in general, could be given in three different forms. The first one is as the value of MIs in a particular point in the phase-space. In this case we can construct a well-defined Cauchy problem. In order to obtain the boundary conditions in this form we need to solve the MIs explicitly in that particular point of the phase-space, either analytically or numerically. If we choose to tackle the problem analytically, one usually choose a point in the phase space in which most of the kinematic invariants vanish, so that the integrals are easier to evaluate. However, a Feynman integral usually develops divergences as we approach such a point, and hence by simply plugging these values into the integrand, we would not obtain the correct asymptotic limit. One possible solution to this problem is given by the method of expansion by regions [34, 35, 36]. Another possibility is to obtain them numerically in the Euclidean region. This can be done, for example, with pySecDec [37], which parametrise the integral and then perform the parametric integration with Monte-Carlo techniques. The problem of this approach is that the numerical precision is limited. Finally, one can use the Auxiliary Mass Flow method, implemented in the Mathematica package AMFlow [16].
There are, then, other two possibilities for fixing the free parameter that appears in the solution. It may happen, indeed, that from independent arguments we know that the integral must be regular in a particular point of the phase-space, usually a pseudo-threshold. In this cases, it might be possible to impose the finiteness of the solution to completely fix the free parameter. A final possibility is to fix the asymptotic behavior of the solution on a threshold. In particular, a general observation we can do about Feynman integrals is that their divergences are no more than logarithmic. In some cases fixing the coefficient of the logarithmic term of the solution might be sufficient to completely determine it. The choice of a method or the other depends on the particular case we are considering.
@@ 4.1 Bottom-to-top approach and solution of equations
$\frac{\partial}{\partial s}\begin{pmatrix}I_{1}^{(\bar{k})}\\ I_{2}^{(\bar{k})}\\ I_{3}^{(\bar{k})}\\ \vdots\\ I_{n}^{(\bar{k})}\end{pmatrix}=\begin{pmatrix}\star&\star&\star&\dots&\star\\ 0&\star&\star&\dots&\star\\ 0&0&\star&\dots&\star\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&\star\\ \end{pmatrix}\begin{pmatrix}I_{1}^{(\bar{k})}\\ I_{2}^{(\bar{k})}\\ I_{3}^{(\bar{k})}\\ \vdots\\ I_{n}^{(\bar{k})}\end{pmatrix}+\begin{pmatrix}\star\\ \star\\ \star\\ \vdots\\ \star\end{pmatrix}.$ (20)
@@ 4.3 Variation of parameters
[…]
Now that we have a particular solution we can construct the general one by summing the homogeneous togheter with the particular one:
@@ 4.4 Example: 1L QED Vertex
$\begin{cases}&r\;c_{0}=-\frac{1}{2}c_{0}\\ &(1+r)\;c_{1}=\frac{1}{8}-\frac{c_{1}}{2}\\ &(2+r)\;c_{2}=\frac{1}{32}\left(1+4c_{1}-16c_{2}\right)\\ &(3+r)\;c_{3}=\frac{1}{128}\left(1+4c_{1}+16c_{2}-64c_{3}\right)\\ &\dots\end{cases}$ (31)
The indicial equation gives $r=-1/2$ and then by recursively solving for all the $c_{i}$ we find:
@@
Finally a comment regarding the convergence of the series centered in $x=0$. In particular, the radius of convergence is given by the distance between the center of the series with the closest singularity. In this case the radius of convergence is $4$, which means that the series converges in the interval $(-4,4)$. This can be seen explicitly in Fig. 1.
Figure 1: Comparison of the series solution for $B_{2}^{(-1)}$ against the exact solution. The series is centered in $x=0$ and the closest singularity is $x=4$. This implies that the series converges in the interval $(-4,4)$, which can be seen from the rapidly growing of the solution around $x=-4$.
@@ 4.5 Coupled Equations
In some practical case might be too complicated, or even not possible if the problem has an elliptic nature, to cast the system of differential equations in the form of Eq. 20. In this section we will discuss how to generalize the methods introduced in the previous sections to system of coupled equations. Let us start by considering the following system of coupled equations
@@
Now we proceed in the same way as the single equation case, that is we substitute it in the homogeneous equation, we expand everything around $x=1$, collect the different powers of $(x-1)$ and, finally, solve recursively the infinite set of algebraic equations for $a_{i}$ and $b_{i}$. The result is:
@@
Firstly, we try to solve the system for $m=0$. If we obtain $n$ linearly independent solutions we are done, otherwise we increase the value of $m$, eventually up to $n-1$.
@@ 4.6 Variation of parameters for systems
[…]
@@
$B_{1}^{(}x)=$ $\frac{59}{8}+\frac{\pi^{2}}{4}+\frac{3}{8}(x-1)+\frac{1}{4}(x-1)^{2}-\frac{1}{3}(x-1)^{3}+\mathcal{O}(x-1)^{4}$
@@ 5.2 Singularities and branch cuts
We saw also in the previous sections, that the solution might exhibit a logarithmic behavior. This complicates the analytic continuation procedure since it makes the solution a multi-valued function. In order to make it single-valued we have to introduce some branch-cuts, thus specifying which is the physical Riemann sheet on which we evaluate the solution. We decided to associate to each logarithmic singularity a branch-cut that starts from the singularity and goes to $\infty$- ∞, parallel to the real axis. While this branch-cut does not modify the region in which the series converges, it shrinks the area in which the solution converges to the desired value, that is the one on the physical Riemann sheet. Looking at left panel of Fig. 3, while the series still converges in $\Gamma_{0}$, only in the blue strip it does to the desired value. As a consequence, for extending the solution we have to choose a path that does not cross the branch cut. This is shown in the right panel of Fig. 3. For reaching $z_{1}$ we cannot go straight along the dashed line, because we would cross the branch cut and end up on the non physical Riemann sheet. Instead, we should follow the path of the solid red line, that avoids $w_{+}$ on the right. For readability of the picture, we have not plotted all the intermediate steps.
@@
Finally, we discuss how to move along a branch cut. There are two different cases which are shown in Fig. 5. If there is not a singularity between the starting and ending point we can move straight since, by definition, we are never crossing a branch cut. If, on the contrary, we have a singularity between the two points we have to avoid moving either in the upper or lower part of the complex plane. This ambiguity is solved by considering Feynman prescription associated to the kinematical variable. A prescription $+i\delta$ implies an horse-shoe path in the upper half of the complex plane, while a $-i\delta$ in the lower one.

### v2

@@ 1 Introduction
The Standard Model (SM) is the theory that classifies all elementary particles and describes their interactions. During the years it has proven itself very successful in explaining and predicting with extreme precisions a big variety of phenomena, spanning several orders of magnitude. The discovery of the existence of the Higgs boson [1, 2], which has been confirmed in 2012 at the Large Hadron Collider (LHC), is one of its greater success. Despite its incredible achievements, however, the SM has some unsolved problems. For example, it does not include gravity in its description of reality, it is not able to explain dark matter and dark energy nor the matter-antimatter asymmetry. The presence of these problems hints that the Standard Model is not yet complete and an extension is needed. Nowadays there exist multiple theories which are suitable candidates for an extension of the SM, namely Grand Unified Theories (GUT), Supersymmetry, String Theory etc., but no clear evidence of physics beyond the SM has been found. Although during the high luminosity run at the LHC we may still be able to make new exciting discoveries, physicists have started to pursue another path to explore the unknown territory of new physics, that is looking for small deviations between theoretical predictions and experimental data. The presence of these small deviations could give, indeed, precious insight on how to extend the SM. For this reason theoretical physicists in the latest years have focused their efforts on obtaining predictions as accurate as possible. These calculations, however, are very far from trivial and in order to achieve a percent, or even sub-percent, level of accuracy one must have control on all the sources of theory uncertainties. These include, for instance, Parton Distributions Functions (PDFs), higher order corrections, computed either in Quantum Chromodynamics (QCD) or in the Electro-Weak (EW) sector, and input scheme. The EW sector of the Standard Model, indeed, depends on three independent parameters selected from precisely measured quantities; this choice affects the organization of radiative corrections, and controlling it is crucial for high-precision predictions [^1] See for example Sec. 5.1 in [3]..
One of the main bottleneck for the inclusion of higher order corrections is the calculation of the so-called Feynman integrals [4, 5]. Indeed, in state-of-the-art problems their number can grow up to thousands or hundred of thousands, and hence a general and algorithmic approach is necessary for their evaluation. Fortunately, not all of them are independent one from the other, and by exploiting integration-by-parts (IBPs) identities [6] one is able to find a set of independent integrals, such that all the others can be expressed in terms of these ones. These integrals are called Master Integrals (MIs). One of the most effective computational tools for evaluating the MIs is the method of differential equations, firstly proposed by Kotikov [7, 8, 9] and later improved by Remiddi and Gehrmann [10, 11, 12, 13]. The idea behind this method is that these integrals are functions of kinematics variables and internal masses, and by differentiating with respect to one of those variables one is able to obtain a first order linear differential equation whose unknowns are the MIs we are interested in. The problem, hence, shifts from integrating to solving a differential equation. The solution to these equations can be written in terms of one-dimensional iterated integrals, which, in many cases, correspond to some known classes of functions such as Harmonic Polylogarithms (HPLs) [14, 15, 16, 17, 18]. However, when increasing the number of loops, external legs or internal masses, an analytical solution in terms of known classes of functions can become extremely difficult to obtain. For this reason, in latest years the series expansion approach is gaining popularity, thanks also to many public implementations: AMFLow [19], DiffExp [20], Line [21] and SeaSyde [22]. Its main idea is to look for a series solution to the differential equations, so that the we are able to evaluate it easily at any point of the phase-space. A general difficulty that arises when calculating corrections at fixed perturbative order is the presence of internal unstable particles such as Ws and Zs. A complete description of resonances in perturbation theory requires a Dyson summation of self-energy insertions. In particular, it has been observed that the pole of the resummed propagator is a gauge-invariant complex quantity. Therefore it is possible to define the renormalized mass of unstable particles as the complex pole of the resummed propagator. This scheme is called complex mass scheme (CMS) [23, 24], and is necessary to correctly describe the kinematic region near the resonance. In order to perform a complete calculation in the CMS we need to be able to evaluate the master integrals also with complex internal masses, and the series expansion approach to the differential equation method can be exploited to this end.
@@ 2.1 Example: two-loop massless triangle
Table 1: Relations between scalar products and denominators for the two-loop massless triangle
@@ 2.2 Integration-by-parts identities
During state-of-the-art calculations, one usually has to compute up to hundred of thousands of FIs. However, not all of them are independent, but there exist linear relations among them. These relations go under the name of Integration-by-parts Identities (IBPs) and they follow from Gauss’s theorem in d-dimensions:
@@
where $v_{\mu}$ can be either a loop or an external momentum. The first step is to write down identities from Eq. 4, for different values of $v_{\mu}$ and for all the $\boldsymbol{\alpha}$ that appear in our calculation. Then we can solve the system by Gauss substitution rule, i.e. considering each equation, expressing one integral in terms of the others and substitute it in the remaining ones. This procedure can be approached in a systematic way by using Laporta algorithm [26]. This is done by assigning a weight to each integral and by systematically expressing integrals of higher weight in terms of others with lower weight. Laporta algorithm is extremely simple and elegant and it has been implemented in several public codes, such as AIR [27], Blade [28], FIRE [29, 30], Kira [31, 32], LiteRed [33], NeatIBP [34] and Reduze [35]. Its algebraic complexity, however, grows very fast with the number of loops, external legs and internal masses. For this reason, for state-of-the-art problem, Laporta algorithm is used in combination with a more advanced technique, the method of finite fields. The idea behind this method is to to solve the system of IBPs numerically multiple times over finite fields and then, at the end, reconstruct the symbolic rational coefficients for the identities of interest by combining the samples together. The advantage of this approach is that the implementation of modular arithmetic in statically typed languages such as C or C++ is fast, exact and suitable for parallelization since each sample is independent from the others. The idea was originally introduced in [36], and later implemented in public packages such as FiniteFlow [37], FireFly [38] and Ratracer [39]. In general, obtaining compact IBPs in a reasonable amount of time is one of the bottleneck of the calculations.
After the implementation of IBPs we are able to express every integral, belonging to the a given integral family, in terms of a finite number of FIs which are called Master Integrals (MIs). These Master Integrals play the role of a basis in the space of Feynman integrals. Note that the term basis may be misleading, because a priori we do not know if this set of MIs is minimal, that is if there exist other relations between them and so if they are truly independent on from the other, anyway this term gives a good idea of what is going on.
@@ 3.2 Example: 1L QED vertex
A second comment regards the number of variables that appear in the system. A common procedure is to introduce adimensional variables, defined as ratios of kinematical invariants. This is done in order to reduce by one the total number of variables in the differential equations, thus simplifying the solution procedure. For example, one could introduce the variable $x=s/m^{2}$, so that the system in Eq. 13 becomes
@@ 3.3 Boundary Conditions
Before diving into how to solve the system of differential equations let us address the problem of finding the boundary conditions. The boundary conditions, in general, could be given in three different forms. The first one is as the value of MIs in a particular point in the phase-space. In this case we can construct a well-defined Cauchy problem. In order to obtain the boundary conditions in this form we need to solve the MIs explicitly in that particular point of the phase-space, either analytically or numerically. If we choose to tackle the problem analytically, one usually chooses a point in the phase space in which most of the kinematic invariants vanish, so that the integrals are easier to evaluate. However, a Feynman integral usually develops divergences as we approach such a point, and hence by simply plugging these values into the integrand, we would not obtain the correct asymptotic limit. One possible solution to this problem is given by the method of expansion by regions [44, 45, 46]. Another possibility is to obtain them numerically in the Euclidean region. This can be done, for example, with pySecDec [47], which parametrises the integral and then perform the parametric integration with Monte-Carlo techniques. The problem of this approach is that the numerical precision is limited. Finally, one can use the Auxiliary Mass Flow method, implemented in the Mathematica package AMFlow [19].
There are, then, other two possibilities for fixing the free parameter that appears in the solution. It may happen, indeed, that from independent arguments we know that the integral must be regular at a particular point of the phase-space, usually a pseudo-threshold. In this cases, it might be possible to impose the finiteness of the solution to completely fix the free parameter. A final possibility is to fix the asymptotic behavior of the solution on a threshold. In particular, a general observation we can do about Feynman integrals is that their divergences are no more than logarithmic. In some cases fixing the coefficient of the logarithmic term of the solution might be sufficient to completely determine it. The choice of a method or the other depends on the particular case we are considering.
@@ 4.1 Bottom-to-top approach and solution of equations
[…]
@@ 4.3 Variation of parameters
$f_{part}(s)=f_{hom}(s)\int_{\tilde{s}}^{s}h(s^{\prime})f_{hom}^{-1}(s^{\prime})ds^{\prime}.$ (27)
Now that we have a particular solution we can construct the general one by summing the homogeneous together with the particular one:
@@ 4.4 Example: 1L QED Vertex
$\begin{cases}&r\;c_{0}=-\frac{1}{2}c_{0}\\ &(1+r)\;c_{1}=\frac{1}{8}-\frac{c_{1}}{2}\\ &(2+r)\;c_{2}=\frac{1}{32}\left(1+4c_{1}-16c_{2}\right)\\ &(3+r)\;c_{3}=\frac{1}{128}\left(1+4c_{1}+16c_{2}-64c_{3}\right)\\$ $\dots\end{cases}$ (31)
The indicial equation gives r=1/2r=-1/2 and then by recursively solving for all the $c_{i}$ we find:
@@
Finally a comment regarding the convergence of the series centered in $x=0$. In particular, the radius of convergence is given by the distance between the center of the series with the closest singularity. In this case the radius of convergence is $4$, which means that the series converges in the interval (4,4)(-4,4). This can be seen explicitly in Fig. 1.
Figure 1: Comparison of the series solution for $B_{2}^{(-1)}$ against the exact solution. The series is centered in $x=0$ and the closest singularity is $x=4$. This implies that the series converges in the interval (4,4)(-4,4), which can be seen from the rapidly growing of the solution around x=4x=-4.
@@ 4.5 Coupled Equations
In some practical case might be too complicated, or even not possible if the problem has an elliptic nature, to cast the system of differential equations in the form of Eq. 20. In this section we will discuss how to generalize the methods introduced in the previous sections to a system of coupled equations. Let us start by considering the following system of coupled equations
@@
Now we proceed in the same way as the single equation case, that is we substitute it in the homogeneous equation, we expand everything around $x=1$, collect the different powers of (x1)(x-1) and, finally, solve recursively the infinite set of algebraic equations for $a_{i}$ and $b_{i}$. The result is:
@@
Firstly, we try to solve the system for $m=0$. If we obtain $n$ linearly independent solutions we are done, otherwise we increase the value of $m$, eventually up to n1n-1.
@@ 4.6 Variation of parameters for systems
$\vec{B}^{part}(x)=\mathbf{M}(x)\int_{x_{0}}^{x}\mathbf{M}^{-1}(x^{\prime})\vec{g}(x^{\prime})dx^{\prime}$ (43)
@@
$B_{1}(x)=$ $\frac{59}{8}+\frac{\pi^{2}}{4}+\frac{3}{8}(x-1)+\frac{1}{4}(x-1)^{2}-\frac{1}{3}(x-1)^{3}+\mathcal{O}(x-1)^{4}$
@@ 5.2 Singularities and branch cuts
We saw also in the previous sections, that the solution might exhibit a logarithmic behavior. This complicates the analytic continuation procedure since it makes the solution a multi-valued function. In order to make it single-valued we have to introduce some branch-cuts, thus specifying which is the physical Riemann sheet on which we evaluate the solution. We decided to associate to each logarithmic singularity a branch-cut that starts from the singularity and goes to -$\infty$, parallel to the real axis. While this branch-cut does not modify the region in which the series converges, it shrinks the area in which the solution converges to the desired value, that is the one on the physical Riemann sheet. Looking at left panel of Fig. 3, while the series still converges in $\Gamma_{0}$, only in the blue strip it does to the desired value. As a consequence, for extending the solution we have to choose a path that does not cross the branch cut. This is shown in the right panel of Fig. 3. For reaching $z_{1}$ we cannot go straight along the dashed line, because we would cross the branch cut and end up on the non physical Riemann sheet. Instead, we should follow the path of the solid red line, that avoids $w_{+}$ on the right. For readability of the picture, we have not plotted all the intermediate steps.
@@
Finally, we discuss how to move along a branch cut. There are two different cases which are shown in Fig. 5. If there is not a singularity between the starting and ending point we can move straight since, by definition, we are never crossing a branch cut. If, on the contrary, we have a singularity between the two points we have to avoid moving either in the upper or lower part of the complex plane. This ambiguity is solved by considering Feynman prescription associated to the kinematical variable. A prescription $i\delta$ implies an horse-shoe path in the upper half of the complex plane, while a $i\delta$ in the lower one.
@@ 6 Conclusions
[…]
Data availability The package SeaSyde is released under the GNU General Public License and it available for download from https://github.com/TommasoArmadillo/SeaSyde.